OpenStudy (lena772):

I'm trying to help a friend out. So the area of a trapezium is 36cm^2 and the expression that was found for the area is x^2+x-6. Do I just put 36^2=x^2+x-6 and solve from there?

OpenStudy (lena772):

I have to solve for x.

OpenStudy (lena772):

@ranga

OpenStudy (lena772):

like then i would put sqrt 36^2 = sqrt x2 +x-6 and get 36 = x+x-6 36=2x-6 42=2x x=21?

OpenStudy (the_fizicx99):

I think so, but you would be multiplying

OpenStudy (the_fizicx99):

I think this is what you tried to do, I haven't taken geo yet but I'm pretty sure you're not suppose to have radicals in your formula

OpenStudy (the_fizicx99):

6 Cannot be simplified, you would have to square it to undo that ._.

OpenStudy (lena772):

this question is freaking stupid.

OpenStudy (the_fizicx99):

If that is the given formula then you would plug in, however for some reason it can't be simplified. Only way I see it is by square rooting both sides to leave 36 alone, but then you would have 36 = x + √x - √6 As far as my knowledge in geometry goes (Pra - Algebra) there isn't suppose to be a radical in that formula. I could be wrong.~

OpenStudy (the_fizicx99):

You would multiply by something ._. wth